OpenStudy (i_ask_alot):

I need help with 7th Grade Math I Medal and Fan the first correct answer!!! Please Help.

OpenStudy (i_ask_alot):

The ratio of girls to boys who participated in the quiz bowl was 7:5. There were 42 girls in the competition. How many boys participated?

OpenStudy (fortytherapper):

Sort everything out. You want to find the number of boys, so let x = boys \[\frac{ Girls }{ Boys} = \frac{ 7 }{ 5 } = \frac{ 42 }{ x}\]

OpenStudy (i_ask_alot):

x= 30?

jhonyy9 (jhonyy9):

how you have got it ?

jhonyy9 (jhonyy9):

cross multiplie and divide both sides by 7 - yes ?

jhonyy9 (jhonyy9):

7x=5*42 x=5*6 x=30

OpenStudy (fortytherapper):

^ Exactly as he said